# Aligning Reasoning LLMs for Materials Discovery with Physics-aware Rejection Sampling

### v1

@@ 1 Introduction
A central goal in materials discovery is to compress the experimental loop by coupling automated experimentation with algorithmic decision making. Within this loop, property prediction is the core module. Accurate models that map composition, structure, and process/recipe variables to target properties (e.g., materials properties or device-level figures of merit) convert combinatorial exploration into tractable optimization and enable closed-loop design via bayesian optimization, provided that the models expose calibrated uncertainty and lightweight physics constraints to keep proposals physically admissible (map_review_2020, ; sdl_review_2024, ; frazier2018_bo_tutorial, ; sabanzagil2025_mfbo, ; jacobs2024_uq_access, ; varivoda2022_uq_benchmark, ). Prior work has largely targeted properties from composition or crystal structure (xie2018_cgcnn, ; chen2019_megnet, ) and, more recently, has explored text- or instruction-conditioned surrogates with LLMs (rubungo2025llm4mat, ); yet scaling these predictors to process-aware, recipe to property tasks at the device level remains challenging (joul22_knowledge_constraints, ; perov23_device_from_recipe, ; nature2024_semiconductor_perspective, ).
@@
Recently, large reasoning models (LRMs)—language models trained and/or reinforced to produce reliable reasoning traces—have shown dominant performance in diverse areas such as math, coding and scientific QA (guo2025deepseek, ; jaech2024openai, ; yang2025qwen3, ). Their step-wise reasoning capabilities are a natural fit for recipe to property prediction, where multi-step physical arguments along the classic chemical composition $\rightarrow$ process $\rightarrow$ micro-structure $\rightarrow$ property chain are essential and well established in integrated computational materials engineering (nrc2008_icme, ). Despite the promise, the leveraging LRMs on property prediction task remains underexplored relative to LLM and naive knowledge extraction (zhao2024_llm_materials_review, ; rubungo2025llm4mat, ). In this paper, we study how to train LRMs that reason effectively about materials recipes and output numerically correct, physically grounded properties.
@@
A prevailing strategy for training LRMs to reason is to use filtered or re-weighted training signals based on the quality of teacher generated traces, complemented by test-time scaling (muennighoff2025s1, ). Concretely, models are fine-tuned on self-generated rationales kept only when they reach correct outcomes (zelikman2022_star, ; yuan2023_rft, ), or on samples ranked by a learned reward/verifier (dong2023_raft, ; cobbe2021_gsm8k, ; zheng2023_llm_judge, ), or they aggregate multiple samples at decoding (wang2022_self_consistency, ). These methods work as standard building blocks for LRMs along with SFT-only pipelines and RL-style post-training (xu2025towards, ; chen2025towards, ).
@@
We argue that training property prediction LRMs from generated reasoning traces requires more sophisticated, physcially grounded rejection sampling. Two characteristics of this task drive the need: (1) High combinatorial design space. The composition, process, structure, property chain creates high-dimensional, multi-mechanism spaces; inverse maps are often non-unique, yielding traces that seem plausible yet are scientifically incorrect (xiang1998_combinatorial, ; takeuchi2004_combinatorial, ; ren2021_inverse_aem, ; liu2018_tandem, ; nsr2022_inverse_generative, ). Therefore, effective learning requires sufficient exploration that searches both the design and reasoning trace space. (2) Physically grounded outputs. Targets are physical quantities whose magnitudes are constrained by physics and even small numeric deviations matter. Filters must therefore enforce admissible ranges and physical constraints from conservation laws and constitutive relations rather than rely solely on binary correctness.
@@
We adopt Qwen3-32B  [^1] https://huggingface.co/Qwen/Qwen3-32B as the backbone model, fine-tuned via supervised fine-tuning (SFT) on internal prompts, with teacher reasoning traces synthesized from Qwen3-235B  [^2] https://huggingface.co/Qwen/Qwen3-235B-A22B (yang2025qwen3, ). We benchmark against various rejection sampling methods under matched token budgets. Empirically, our method achieves the highest overall accuracy and calibration, while also delivering superior compute efficiency compared to existing baselines.
@@
  • We formulate recipe to property prediction as a reasoning task where physics-aware verification is essential.
@@ 2 Related Work
Recent advances in large language models (LLMs) have demonstrated strong generalization capabilities in materials design, drawing on interdisciplinary knowledge from chemistry, physics, and engineering miret2024llms ; jia2024llmatdesign . Beyond general-purpose LLMs, domain-specific models such as MatSciBERT gupta2022matscibert , MatBERT wan2024tokens , and MELT kim2024melt have been trained on large-scale materials science corpora, successfully capturing fundamental concepts that link structure, properties, processes, and performance. A key step in the materials design pipeline is accurate property prediction from symbolic representations, which serves as a surrogate for expensive experiments and enables rapid candidate screening. Leveraging their ability to process unstructured scientific data, recent studies have applied LLMs to this task without the need for elaborate feature engineering. For example, LLM-Prop niyongabo2025llm employs LLMs to predict crystalline material properties, while Li et al.li2025hybrid integrate LLMs with graph neural networks for improved prediction accuracy. In the context of quantum-dot materials, Choi et al.choi2025llm developed LLM-based synthesis protocol generation and property prediction models, fine-tuned on proprietary synthesis datasets. LLMs have also been explored as surrogate models in optimization frameworks. LLAMBO liularge utilizes the exploration capability of LLMs within Bayesian optimization, and BOPRO agarwal2025searching incorporates LLM-based search strategies that exploit evolving uncertainty estimates to propose promising candidates in each iteration, thereby accelerating the discovery of globally optimal solutions.
@@
The paradigm of next-token prediction has undergone a significant shift with the introduction of "thought" concept—a sequence of intermediate steps representing a model’s internal reasoning process jaech2024openai ; guo2025deepseek ; yang2025qwen3 ; muennighoff2025s1 . This innovative approach enables LLMs to mimic complex human reasoning, such as reflective thinking and tree search. Chain-of-Thought (CoT) prompting wei2022chain initially demonstrated that few-shot rationales could unlock sophisticated reasoning in LLMs, while Self-Consistency wangself further enhanced reliability by marginalizing over multiple reasoning paths. Tree-of-Thoughts (ToT) yao2023tree later reframed inference as a search over partial thought sequences, incorporating look-ahead and backtracking. Beyond prompting, process supervision has emerged as a powerful technique, training step-level verifiers or reward models to guide the reasoning process. This approach has been shown to outperform models trained with outcome-only labels, particularly in mathematical reasoning tasks lightman2023let . More recently, GPT-o1 jaech2024openai , Qwen3 yang2025qwen3 , and DeepSeek-R1 guo2025deepseek have popularized modern Large Reasoning Models (LRMs) by integrating long-form thinking with process supervision, RL-based post-training, and test-time scaling. Building on these foundational advances, our work adapts these reasoning mechanisms to the domain of materials design, specifically targeting physically grounded recipe to property prediction.
@@
Rejection sampling is widely recognized as an effective data-filtering method that promotes higher-quality supervision. In the context of RLHF and preference-optimization, it narrows multiple generated outputs per prompt to only the high quality responses, as determined by a reward model during post-training adjustments. For example, RAFTdongraft aligns generative models efficiently by using a reward model and abundant candidate samples; it discards outputs demonstrating undesired behaviors and fine-tunes the model solely on the selected high-quality subset. Building on this, Reinforce-Rejxiong2025minimalist proposes a minimalist policy-gradient extension that filters out both entirely incorrect and entirely correct samples, enhancing stability and efficiency. In reasoning tasks, STAR-like modelszelikman2022star ; hosseini2024v ; koh2025adastar eliminate expensive human annotations by employing a self-taught reasoning loop—generating chain-of-thought traces, self-verifying correctness, and fine-tuning only on reliable examples. Additionally, Rejection Sampling Fine-Tuning (RFT)yuan2024scaling enhances mathematical reasoning by incorporating model-generated reasoning traces filtered for correctness into training.
@@ 3.1 Task Definition
Quantum dot light-emitting diodes (QD-LEDs) are electroluminescent devices that use colloidal semiconductor quantum dots as the emissive layer, offering narrowband spectra and composition tunable color (Shirasaki2013, ; Li2024, ). In practice, closed-loop materials design must optimize figures of merit measured at the device level—e.g., peak external quantum efficiency and operational stability—because these metrics ultimately determine application viability for emissive displays and lighting (e.g., televisions, monitors) (Bang2021, ; Huang2020, ; Li2024, ). We therefore frame inputs as complete device recipes: a multi-layer stack typically consist of anode/ITO, hole-injection layer (HIL), hole-transport layer (HTL), quantum-dot emitting layer (EML), electron-transport layer (ETL), electron-injection layer/cathode with per layer materials and process parameters. Each layer records identifiers (material, formulation), geometric parameters (e.g., thickness), and process variables (e.g., solution concentration, spin profile, bake/anneal temperature and duration, atmosphere), along with post-process steps (e.g., UV–ozone, plasma, solvent rinse). See Figure 1 for a recipe example.
@@ 3.2 PaRS: Physics-aware Rejection Sampling
As $\mathrm{EQE}_{\max}$ is reported as a percentage, Eq. (2) enforces range consistency. Unlike categorical correctness filters (zelikman2022_star, ; yuan2023_rft, ; dong2023_raft, ; wang2022_self_consistency, ; muennighoff2025s1, ), Eq. (3) uses a continuous error against wet-lab ground truth, yielding richer learning signals.
@@
With external quantum efficiency (EQE) as the prediction target, we can define an empirical upper bound $U(x)$ as follows. EQE is commonly factorized as $\mathrm{EQE}\;=\;\eta_{\mathrm{out}}\cdot\eta_{\mathrm{rad}}\cdot\gamma$ with $\eta_{\mathrm{out}}\leq 1$ and $\gamma\leq 1$ by definition. In regimes where solid-state PLQY limits the radiative yield, the conservative relation $\eta_{\mathrm{rad}}\leq\mathrm{PLQY}$ holds, implying $\mathrm{EQE}\leq\mathrm{PLQY}$ (Shirasaki2013NatPhotonics, ; GatherReineke2015JPE, ; OE2014_QLED_dipole, ). PLQY (photoluminescence quantum yield) denotes the ratio of emitted to absorbed photons; we use the solid-state PLQY of the EML under device-relevant conditions. We therefore define an empirical, recipe-specific upper bound as $U(x){=}U_{\mathrm{PL}}(x)$, where $U_{\mathrm{PL}}(x)$ is instantiated by the highest measured film photoluminescence quantum yield of the emissive layer in the recipe $x$. We add this upper bound to the acceptance gates:
@@ 4.1 Baselines
    Self-consistency aggregation: select the trace whose final answer is closest to the median across all $K$ answers (wang2022_self_consistency, ).
@@
    LLM-as-a-judge: score all traces with a larger judge model (DeepSeek-R1  [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-0528) and select the top-ranked trace (zheng2023judging, ). Refer Appendix. A.2 for details of the LLM-as-a-Judge prompting.
@@
    Multi-sampling: retain all $K$ traces as supervision (guha2025openthoughts, ).
@@
Table 2: QD-LED property prediction with Qwen3-32B trained on teacher traces selected by each method from Qwen3-235B. Lower MAE is better; higher $R^{2}$ and Spearman $\rho$ are better. Viol.% is computed on test predictions without post hoc clipping to $U(x)$. # train is the number of supervision traces used for SFT: multi-sampling retains all K=12K{=}12 traces per example (thus $12\times$ larger), whereas our method keeps only candidates that pass our gates, resulting in fewer traces.
Training data (prompt + reasoning trace) # train MAE (pp) $\boldsymbol{R^{2}}$ Spearman $\boldsymbol{\rho}$ Viol. (%) no rejection sampling 10 00010\,000 $2.001$ $0.376$ $0.607$ $35.8$ random sampling 10 00010\,000 $1.961$ $0.358$ $0.621$ 35.2 longest sampling 10 00010\,000 $1.942$ $0.375$ $0.614$ $35.3$ self-consistency sampling 10 00010\,000 $1.933$ $0.377$ $0.629$ $36.8$ LLM-as-a-Judge 10 00010\,000 1.889 0.408 0.667 $35.4$ multi-sampling 120 000120\,000 $1.984$ $0.335$ $0.632$ $36.6$ PaRS (ours) $8000$ $1.808$ $0.424$ $0.705$ $27.7$
[…]
Figure 4: Compute–accuracy frontier for rejection sampling methods. Our approach achieves the lowest student MAE with substantially fewer tokens, forming the empirical Pareto front. The x-axis shows required tokens per example and the y-axis shows student MAE. See Appendix A.3 for details.

### v2

@@ 1 Introduction
A central goal in materials discovery is to compress the experimental loop by coupling automated experimentation with algorithmic decision making. Within this loop, property prediction is the core module. Accurate models that map composition, structure, and process/recipe variables to target properties (e.g., materials properties or device-level figures of merit) convert combinatorial exploration into tractable optimization and enable closed-loop design via bayesian optimization, provided that the models expose calibrated uncertainty and lightweight physics constraints to keep proposals physically admissible  (Stach et al., 2022; Flores-Leonar et al., 2024; Frazier, 2018; Sabanzagil et al., 2025; Jacobs et al., 2024; Varivoda et al., 2022). Prior work has largely targeted properties from composition or crystal structure  (Xie & Grossman, 2018; Chen et al., 2019) and, more recently, has explored text- or instruction-conditioned surrogates with LLMs  (Ndayishimiye et al., 2025); yet scaling these predictors to process-aware, recipe to property tasks at the device level remains challenging  (Liu et al., 2022; Lu et al., 2023; Xie et al., 2024).
@@
Recently, large reasoning models (LRMs)—language models trained and/or reinforced to produce reliable reasoning traces—have shown dominant performance in diverse areas such as math, coding and scientific QA  (Guo et al., 2025; Jaech et al., 2024; Yang et al., 2025). Their step-wise reasoning capabilities are a natural fit for recipe to property prediction, where multi-step physical arguments along the classic chemical composition $\rightarrow$ process $\rightarrow$ micro-structure $\rightarrow$ property chain are essential and well established in integrated computational materials engineering  (National Research Council, 2008). Despite the promise, the leveraging LRMs on property prediction task remains underexplored relative to LLM and naive knowledge extraction  (Zhao et al., 2024; Ndayishimiye et al., 2025). In this paper, we study how to train LRMs that reason effectively about materials recipes and output numerically correct, physically grounded properties.
@@
A prevailing strategy for training LRMs to reason is to use filtered or re-weighted training signals based on the quality of teacher generated traces, complemented by test-time scaling  (Muennighoff et al., 2025). Concretely, models are fine-tuned on self-generated rationales kept only when they reach correct outcomes  (Zelikman et al., 2022b; Yuan et al., 2023), or on samples ranked by a learned reward/verifier  (Dong et al., 2023; Cobbe et al., 2021; Zheng et al., 2023b), or they aggregate multiple samples at decoding  (Wang et al., 2022). These methods work as standard building blocks for LRMs along with SFT-only pipelines and RL-style post-training  (Xu et al., 2025; Chen et al., 2025).
@@
We argue that training property prediction LRMs from generated reasoning traces requires more sophisticated, physically grounded rejection sampling. Two characteristics of this task drive the need: (1) High combinatorial design space. The composition, process, structure, property chain creates high-dimensional, multi-mechanism spaces; inverse maps are often non-unique, yielding traces that seem plausible yet are scientifically incorrect  (Xiang et al., 1995; Takeuchi et al., 2002; Ren et al., 2021; Liu et al., 2018; Yang et al., 2022). Therefore, effective learning requires sufficient exploration that searches both the design and reasoning trace space. (2) Physically grounded outputs. Targets are physical quantities whose magnitudes are constrained by physics and even small numeric deviations matter. Filters must therefore enforce admissible ranges and physical constraints from conservation laws and constitutive relations rather than rely solely on binary correctness.
@@
We adopt Qwen3-32B https://huggingface.co/Qwen/Qwen3-32B as the backbone model, fine-tuned via supervised fine-tuning (SFT) on internal prompts, with teacher reasoning traces synthesized from Qwen3-235B https://huggingface.co/Qwen/Qwen3-235B-A22B  (Yang et al., 2025). We benchmark against various rejection sampling methods under matched token budgets. Empirically, our method achieves the highest overall accuracy and calibration, while also delivering superior compute efficiency compared to existing baselines.
@@
  • We formulate recipe to property prediction as a reasoning task with LRMs where physics-aware verification is essential.
@@ 2 Related Work
Recent advances in large language models (LLMs) have demonstrated strong generalization capabilities in materials design, drawing on interdisciplinary knowledge from chemistry, physics, and engineering  (Miret & Krishnan, ; Jia et al., 2024). Beyond general-purpose LLMs, domain-specific models such as MatSciBERT  (Gupta et al., 2022), MatBERT  (Wan et al., 2024), and MELT  (Kim et al., 2024) have been trained on large-scale materials science corpora, successfully capturing fundamental concepts that link structure, properties, processes, and performance. A key step in the materials design pipeline is accurate property prediction from symbolic representations, which serves as a surrogate for expensive experiments and enables rapid candidate screening. Leveraging their ability to process unstructured scientific data, recent studies have applied LLMs to this task without the need for elaborate feature engineering. For example, LLM-Prop  (Niyongabo Rubungo et al., 2025) employs LLMs to predict crystalline material properties, while Li et al. (Li et al., 2025) integrate LLMs with graph neural networks for improved prediction accuracy. In the context of quantum-dot materials, Choi et al. (Choi et al., 2025) developed LLM-based synthesis protocol generation and property prediction models, fine-tuned on proprietary synthesis datasets. LLMs have also been explored as surrogate models in optimization frameworks. LLAMBO  (Liu et al., ) utilizes the exploration capability of LLMs within Bayesian optimization, and BOPRO  (Agarwal et al., 2025) incorporates LLM-based search strategies that exploit evolving uncertainty estimates to propose promising candidates in each iteration, thereby accelerating the discovery of globally optimal solutions.
@@
The paradigm of next-token prediction has undergone a significant shift with the introduction of "thought" concept—a sequence of intermediate steps representing a model’s internal reasoning process  (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025; Muennighoff et al., 2025). This innovative approach enables LLMs to mimic complex human reasoning, such as reflective thinking and tree search. Chain-of-Thought (CoT) prompting  (Wei et al., 2022) initially demonstrated that few-shot rationales could unlock sophisticated reasoning in LLMs, while Self-Consistency  (Wang et al., ) further enhanced reliability by marginalizing over multiple reasoning paths. Tree-of-Thoughts (ToT)  (Yao et al., 2023) later reframed inference as a search over partial thought sequences, incorporating look-ahead and backtracking. Beyond prompting, process supervision has emerged as a powerful technique, training step-level verifiers or reward models to guide the reasoning process. This approach has been shown to outperform models trained with outcome-only labels, particularly in mathematical reasoning tasks  (Lightman et al., 2023). More recently, GPT-o1  (Jaech et al., 2024), Qwen3  (Yang et al., 2025), and DeepSeek-R1  (Guo et al., 2025) have popularized modern Large Reasoning Models (LRMs) by integrating long-form thinking with process supervision, RL-based post-training, and test-time scaling. Building on these foundational advances, our work adapts these reasoning mechanisms to the domain of materials design, specifically targeting physically grounded recipe to property prediction.
@@
Rejection sampling is widely recognized as an effective data-filtering method that promotes higher-quality supervision. In the context of RLHF and preference-optimization, it narrows multiple generated outputs per prompt to only the high quality responses, as determined by a reward model during post-training adjustments. For example, RAFT (Dong et al., ) aligns generative models efficiently by using a reward model and abundant candidate samples; it discards outputs demonstrating undesired behaviors and fine-tunes the model solely on the selected high-quality subset. Building on this, Reinforce-Rej (Xiong et al., 2025) proposes a minimalist policy-gradient extension that filters out both entirely incorrect and entirely correct samples, enhancing stability and efficiency. In reasoning tasks, STAR-like models (Zelikman et al., 2022a; Hosseini et al., 2024; Koh et al., 2025) eliminate expensive human annotations by employing a self-taught reasoning loop—generating chain-of-thought traces, self-verifying correctness, and fine-tuning only on reliable examples. Additionally, Rejection Sampling Fine-Tuning (RFT) (Yuan et al., 2024) enhances mathematical reasoning by incorporating model-generated reasoning traces filtered for correctness into training.
@@ 3.1 Task Definition
Quantum dot light-emitting diodes (QD-LEDs) are electroluminescent devices that use colloidal semiconductor quantum dots as the emissive layer, offering narrowband spectra and composition tunable color  (Shirasaki et al., 2013a; Li et al., 2024). In practice, closed-loop materials design must optimize figures of merit measured at the device level—e.g., peak external quantum efficiency and operational stability—because these metrics ultimately determine application viability for emissive displays and lighting (e.g., televisions, monitors)  (Bang et al., 2021; Huang et al., 2020; Li et al., 2024). We therefore frame inputs as complete device recipes: a multi-layer stack typically consist of anode/ITO, hole-injection layer (HIL), hole-transport layer (HTL), quantum-dot emitting layer (EML), electron-transport layer (ETL), electron-injection layer/cathode with per layer materials and process parameters. Each layer records identifiers (material, formulation), geometric parameters (e.g., thickness), and process variables (e.g., solution concentration, spin profile, bake/anneal temperature and duration, atmosphere), along with post-process steps (e.g., UV–ozone, plasma, solvent rinse). See Figure 1 for a recipe example.
@@ 3.2 PaRS: Physics-aware Rejection Sampling
As $\mathrm{EQE}_{\max}$ is reported as a percentage, Eq. (2) enforces range consistency. Unlike categorical correctness filters  (Zelikman et al., 2022b; Yuan et al., 2023; Dong et al., 2023; Wang et al., 2022; Muennighoff et al., 2025), Eq. (3) uses a continuous error against wet-lab ground truth, yielding richer learning signals.
@@
With external quantum efficiency (EQE) as the prediction target, we can define an empirical upper bound $U(x)$ as follows. EQE is commonly factorized as $\mathrm{EQE}\;=\;\eta_{\mathrm{out}}\cdot\eta_{\mathrm{rad}}\cdot\gamma$ with $\eta_{\mathrm{out}}\leq 1$ and $\gamma\leq 1$ by definition. In regimes where solid-state PLQY limits the radiative yield, the conservative relation $\eta_{\mathrm{rad}}\leq\mathrm{PLQY}$ holds, implying $\mathrm{EQE}\leq\mathrm{PLQY}$  (Shirasaki et al., 2013b; Gather & Reineke, 2015; OE2, 2014). PLQY (photoluminescence quantum yield) denotes the ratio of emitted to absorbed photons; we use the solid-state PLQY of the EML under device-relevant conditions. We therefore define an empirical, recipe-specific upper bound as $U(x){=}U_{\mathrm{PL}}(x)$, where $U_{\mathrm{PL}}(x)$ is instantiated by the highest measured film photoluminescence quantum yield of the emissive layer in the recipe $x$. We add this upper bound to the acceptance gates:
@@ 4.1 Baselines
    Self-consistency aggregation: select the trace whose final answer is closest to the median across all $K$ answers  (Wang et al., 2022).
@@
    LLM-as-a-judge: score all traces with a larger judge model (DeepSeek-R1 §§§https://huggingface.co/deepseek-ai/DeepSeek-R1-0528) and select the top-ranked trace  (Zheng et al., 2023a). Refer Appendix. A.2 for details of the LLM-as-a-Judge prompting.
@@
    Multi-sampling: retain all $K$ traces as supervision  (Guha et al., 2025).
@@
$\dagger$ We drop the around 20% of sample that not passed our acceptance logic in Sec. 3.2 thus $0.8$ traces kept per prompt on average.
@@
Table 2: QD-LED property prediction with Qwen3-32B trained on teacher traces selected by each method from Qwen3-235B. Lower MAE is better; higher $R^{2}$ and Spearman $\rho$ are better. Viol.% is computed on test predictions without post hoc clipping to $U(x)$. # train is the number of supervision traces used for SFT: multi-sampling retains all K=12K{=}12 traces per example ($12\times$ number of train prompts), whereas our method keeps only candidates that pass our gates, resulting in fewer traces.
Training data (prompt + reasoning trace) # train MAE $\boldsymbol{R^{2}}$ Spearman $\boldsymbol{\rho}$ Viol. (%) No sampling 10 00010\,000 $2.001$ $0.376$ $0.607$ $35.8$ Random sampling 10 00010\,000 $1.961$ $0.358$ $0.621$ 35.2 Longest trace 10 00010\,000 $1.942$ $0.375$ $0.614$ $35.3$ Self-consistency 10 00010\,000 $1.933$ $0.377$ $0.629$ $36.8$ LLM-as-a-Judge 10 00010\,000 1.889 0.408 0.667 $35.4$ Multi-sampling 120 000120\,000 $1.984$ $0.335$ $0.632$ $36.6$ PaRS (ours) $7980$ $1.808$ $0.424$ $0.705$ $27.7$
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.61.81.8222.22.22.42.42.62.600.50.5111.51.5222.52.5empirical pareto-frontAvg. tokens to generate trace ($10^{5}$)MAEPaRS (ours)Self-consistencyRandomLLM-as-a-Judge
Figure 4: Compute–accuracy frontier for rejection sampling methods. Our approach achieves the lowest teacher MAE with substantially fewer required tokens, forming the empirical Pareto front. The x-axis shows average required tokens for generating reasoning trace per prompt and the y-axis shows teacher MAE. See Appendix A.3 for details.